OpenStudy (konradzuse):

In 1965, about 44% of the U.S. adult population had never smoked cigarettes. A national health survey of 1205 U.S. adults (presumably selected randomly) during 2006 revealed that 615 had never smoked cigarettes. Reference: Ref 18-5 Suppose you wished to test whether there has been a change since 1965 in the proportion of U.S. adults that have never smoked cigarettes. Which of the following are the appropriate hypotheses? A. H0: p = 0.44, Ha: p > 0.44 B. H0: p = 0.51, Ha: p 0.51 C. H0: p = 0.44, Ha: p 0.44 D. H0: = 0.44, Ha: 0.44 The P-value of the test of hypotheses is A. greater than 0.10. B. between 0.05 and 0.10. C. between .01 and 0.05. D. below 0.01.

OpenStudy (konradzuse):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what did you get for the first part

OpenStudy (konradzuse):

I forgot to mention that there was a misprint when I pasted.. B is not equal, same with C, and then there is a p ^ for both of D's answers... I think it's B....

jimthompson5910 (jim_thompson5910):

D is out because you're not testing p^, you're testing p

jimthompson5910 (jim_thompson5910):

how are you getting 51% ? the given percentage is 44%

jimthompson5910 (jim_thompson5910):

oh i see, you divided 615/1205 = 0.51, but that's p^ and NOT p

OpenStudy (konradzuse):

I meant C?

jimthompson5910 (jim_thompson5910):

you did eh? lol yeah C is correct

OpenStudy (konradzuse):

Haha I glanced over them didn't realize that there were 2 of the same haha... I wanted the one that was not equal :D.

OpenStudy (konradzuse):

Would the next part be B?

jimthompson5910 (jim_thompson5910):

you first need to find the test statistic

OpenStudy (konradzuse):

hmm....

jimthompson5910 (jim_thompson5910):

use the formula z = (m - p)/(sqrt(pq/n)) where m is the test statistic and q = 1-p z = (m - p)/(sqrt(pq/n)) z = (0.51037 - 0.44)/(sqrt(0.44*0.56/1205)) z = 4.92108 So the test statistic is z = 4.92108

jimthompson5910 (jim_thompson5910):

now you need to find P(z > 4.92108) and double this result

jimthompson5910 (jim_thompson5910):

you can get away without even using a calculator though because notice how far z = 4.92108 is from z = 0

OpenStudy (konradzuse):

mhm,.

jimthompson5910 (jim_thompson5910):

so what's the p value

OpenStudy (konradzuse):

not b hmm? A?

jimthompson5910 (jim_thompson5910):

what did you get for the p-value, which numeric value did you get

OpenStudy (konradzuse):

not sure....

jimthompson5910 (jim_thompson5910):

calculate P(z > 4.92108), what do you get

OpenStudy (konradzuse):

How do we calculate that again without the table?

jimthompson5910 (jim_thompson5910):

type normalcdf(4.92108, 10) into wolfram alpha

jimthompson5910 (jim_thompson5910):

or you can use a TI calculator to do the same thing

OpenStudy (konradzuse):

hm.... Now the le end poin is almos 5, but the oher bnumbers are super tiny

jimthompson5910 (jim_thompson5910):

so that means what

OpenStudy (konradzuse):

A or D.

jimthompson5910 (jim_thompson5910):

why A

OpenStudy (konradzuse):

Well like I said I wasn't sre if it was the 5 or tiny #'s.. So D then :)

jimthompson5910 (jim_thompson5910):

well the p value is very very small, so it has to be less than 0.01

jimthompson5910 (jim_thompson5910):

the p value is the 4th value down

jimthompson5910 (jim_thompson5910):

well you double it first, but still, it's very small